What are the net present value and internal rates-of-return of the irrigation system if the required-rate-of-return is 10%? Is the investment acceptable? knowing that irrigation system costs $25,000 and has returns of: $12,500 in year one, $10,000 in year two, $7,500 in year three, and $5,000 in year four.
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=12500/1.1+10,000/1.1^2+7500/1.1^3+5000/1.1^4

=28678.03

NPV=Present value of inflows-Present value of outflows

=28678.03-25000

=$3678.03(Approx).

Let irr be x%
At irr,present value of inflows=present value of outflows.

25000=12500/1.0x+10,000/1.0x^2+7500/1.0x^3+5000/1.0x^4

Hence x=irr=17.81%(Approx.